The sum of two numbers is seventeen. Twice the smaller number is four more than the larger number.
Answer by tsewaitung(11)

let the smaller number be x and the larger be y.
from the equation, we have
x+y=17------(1)
2x-4=y------(2)
by substituting (2) into (1), we have
x+2x-4=17
3x=21
x=7
by substituting x=7 into (2), we have
y=2(7)-4
y=14-4
y=10